OpenStudy (anonymous):

The astronomy club wants to buy five different sizes of eyepieces for a new telescope. The 6-, 12-, and 20-millimeter lenses cost $42.95 each, and the 32- and 40-millimeter lenses cost $57.96 each. Estimate the total cost of the lenses to the nearest whole dollar. $345 $166 $245

OpenStudy (anonymous):

What seems to be the confusion? There are 3 lenses that cost $42.95 each, so (42.92 * 3) = $128.85 There are 2 different lenses that cost $57.96 each (57.96 * 2) = $115.92 So add them together, that is your total, and round to nearest dollar. 128.85 + 115.92 = 244.77 The nearest whole dollar, is 245. So the total cost would be $245 to the nearest dollar.

OpenStudy (anonymous):

im 7 but thank you very muich!!!!